Use the compound Al(NO3)3 to answer the following questions:

1. What is the molar mass of this compound? Show your work.

2. How many moles of this compound are found in 5 grams of the compound? Show your work.

3. How many molecules is this? Show your work.
Answer #1

1.

Molar mass = 1×27+3×14+9×16 = 213.0 g/mol

2. Moles = mass/molar mass

Mole = 5.0/213.0 => 0.0235 mol

3. Number of molecues = moles ×NA

=> 0.0235×6.022×1023

= 1.41×1022 molecues